# Prompt-Counterfactual Explanations for Generative AI System Behavior

### v1

@@ 4.1 Explanation algorithm
In the initial scoring phase, the PCE-1 algorithm evaluates the influence of candidate explanations (elements in the input prompt) on a specific output characteristic such as toxicity or sentiment. First, the original prompt $S$ is passed to a generative AI system $G$ to produce a number of outputs (with the number equal to num_samplesnum\_samples), presuming G is non-deterministic. These generated outputs are evaluated using downstream classifier $C_{m}$, which returns a score $p$ indicating how strongly the output exhibits the focal target property. As explained above, how this score is calculated depends on the downstream tasks. Next, if this score exceeds the predefined thresholdthreshold, the algorithm proceeds to an element-level sensitivity analysis for that prompt. For each element $e_{i}$ in the prompt (with in total $n$ elements), a new version of the prompt is created with that element replaced (or masked) ($S\setminus e_{i}$). The modified prompt is fed again into the generative model, and the resulting outputs are scored by the classifier. These element-level scores are recorded in a dictionary, capturing how the removal of each element affects the output’s classification score.
@@
This explanation algorithm is intended as a first proof of concept for generating PCEs for a non-deterministic, generative AI system. Our default search strategy is an element-level textual masking approach, where we iteratively replace (“mask”) elements from the input with underscores ($_{}^{\prime}\_^{\prime}$) to test their causal effect on the classifier’s output. However, more sophisticated techniques could be explored, such as using system-specific masking tokens. As an alternative to masking, replacing input segments with alternative text may yield more natural and actionable counterfactuals (e.g, kim-etal-2020-interpretation). In settings with very large input prompts, another alternative would be grouping phrases or semantically related elements and evaluating their collective influence as an initial filtering step could substantially improve the efficiency of the search process.
This approach could be applied to any large language model. For the experiments in this paper, we apply it to LLaMA 3.1–8B (grattafiori2024llama) (hereafter LLaMA) and OLMo-2-0425-1B (walsh20252) (hereafter OLMo). [^14] https://huggingface.co/meta-llama/Llama-3.1-8B, [^15] https://huggingface.co/allenai/OLMo-2-0425-1B We use the following parameter settings: num_samplesnum\_samples = 10 (except for toxicity, where we use num_samplesnum\_samples = 100) and num_explanationsnum\_explanations = 5, and time_limittime\_limit = 60 seconds. The returned numbers of explanations for each domain and model are listed in Table 1.
@@ 5 Case studies
We illustrate the production of PCEs for generative (LLM-based) systems with three case studies: [^16] Similarly to the prior work of fernandez_explaining. detection of political leaning, toxicity prediction, and sentiment classification. Each case study shows the explanations produced with case-specific settings for the algorithm parameters, and discusses case-specific implications. The case studies also demonstrate the versatility of counterfactual explanations across tasks with very different data distributions: in political leaning detection, the distribution is bimodal; in toxicity detection, the signal is concentrated in the relatively few toxic instances; and in sentiment analysis, the focus is on the long tail.
@@ 5.1 Case Study 1: Detection of political leaning
LLM-based AI systems have been criticized for producing biased outputs (gallegos2024bias; ho2025gender; goethals2026fairness). One type of bias that is often examined is when the generated outputs systematically lean toward one side of the political spectrum (motoki2024more; rettenberger2025assessing; anthropic2025political). Stakeholders investing in, building, and integrating AI systems are interested in (a) whether their AI systems’ generations do in fact lean one direction or the other politically, (b) if so, are there specific things in the prompts that cause the systems to generate such biased outputs? And (c) what they might do about it. Prior work has created classifiers to address (a), which we discuss below. This paper presents a solution to (b). A full answer to part (c) is beyond the scope of this paper (and is very complex); we will provide some ideas for (c) toward the end of this section. [^17] OpenAI recently reported on their desire that their LLM-driven AI systems not exhibit political bias, as well as what they are doing about it, stating clearly that “ChatGPT shouldn’t have political bias in any direction.” They estimate that in live traffic “less than 0.01% of all ChatGPT responses show any signs of political bias.” We note that this still would result in a non-negligible number of biased responses, given the tremendous volume of queries to ChatGPT. https://openai.com/index/defining-and-evaluating-political-bias-in-llms/. Similarly, Anthropic announced in a blog post that they are reducing and actively monitoring political bias (anthropic2025political).
@@
Generating news articles from headlines is a well-established method for evaluating political bias in language models (bang2024measuring). To assess the political leaning of the generated outputs, we rely on PoliticalBiasBERT, a classifier designed to predict the dominant political ideology of news content that is widely used (bucket_bias2023; baly2020we). This model produces as output one of three categorical labels—left, center, or right. As noted above, we apply PoliticalBiasBERT not to the headlines, but to the generated output—in line with the architecture in Figure 1. Thus the headline is the prompt and the news article is the generated output, which aligns with many uses of LLM-based systems, taking a relatively short prompt and producing a significantly longer output.
@@
As discussed in the literature on counterfactual explanation (martens2014explaining; wachter2017counterfactual; verma2024counterfactual), there are many goals for producing such explanations. One common goal is to help managers and developers understand the reasons for undesirable system behavior. These results illustrate how PCEs can inform these stakeholders about the elements of the input prompts that led to the political leaning.
@@
We see from the figure that there are indeed words that consistently produce right-leaning output. Such an analysis could be helpful to data scientists, developers, and domain experts working to improve the AI model; a common argument for including counterfactual explanations in the model-development toolkit (martens2014explaining; abid2022meaningfully; gan2021automated; yousefzadeh2019debugging). Specifically, the development process can focus on understanding why the presence of these particular words in prompts leads to the biased output, and gives a focused target for reducing the bias.
@@
Deciding what to do based on what counterfactual explanations reveal depends not only on the problem setting but also on the stakeholder in question (langer2021we). A top-notch AI engineer would have more options at her disposal than a business developer working on an LLM-driven solution. One potential broad application of PCEs for generative outputs is prompt engineering. Knowing specifically what caused a desired or undesired output characteristic could help to engineer future prompts that produce or avoid desired or undesired generations. A comprehensive investigation of prompt engineering driven by PCEs is well beyond the scope of this paper (and would be a significant contribution as future work). As a follow-up for this case study, we demonstrate the potential as follows.
The generalizability of the content of the explanations suggests that the PCEs could inform prompt engineering strategies for avoiding the focal characteristic (here, right-leaning) in the LLM generations. To illustrate, we conducted the following comparison. We (a) again selected the 100 headlines that contained the highest frequency of the top 20 politically salient words, and (b) created a corresponding set of headlines in which we replaced these top 20 words with synonyms (e.g., replacing “lawsuit” with “legal case” and “Netanyahu” with “Israeli prime minister”). [^19] We asked ChatGPT to provide synonyms for each of the words. The results are shown in Table 4.
@@ 5.2 Case Study 2: Toxicity prediction
LLMs carry a well-documented risk of producing toxic or otherwise objectionable content. This issue to a large extent arises from the nature of their training data—large-scale internet corpora that inevitably include hate speech, stereotypes, and offensive language (abid2021persistent; bender2021dangers; gehman2020realtoxicityprompts). As a result, even when prompted with seemingly neutral inputs, LLMs can generate responses that violate corporate policies and social or content-safety norms. While these risks remain a concern, they are substantially mitigated through extensive post-training interventions implemented by the providers. Nonetheless, although rare, the systems still sometimes generate toxic outputs. The consequences of such outputs range from diminished trust in AI systems to reputational and legal risks for the organizations deploying them (weidinger2022taxonomy). Thus both LLM providers and organizations integrating them have an interest in understanding and avoiding toxic generations.
To study the relationship between prompts and toxic generated outputs, we use the RealToxicityPrompts dataset (gehman2020realtoxicityprompts), which contains a mix of prompts that either are neutral or were known to induce toxic outputs under some conditions. [^20] https://huggingface.co/datasets/allenai/real-toxicity-prompts Although the vast majority of prompts are unlikely to yield toxic responses from current LLMs, we observe that toxicity indeed still emerges. To determine whether a given output is toxic, we use the RoBERTa-based toxicity classifier developed by logacheva-etal-2022-paradetox, which outputs a binary label for toxicity. [^21] https://huggingface.co/s-nlp/roberta_toxicity_classifier This case study examines the setting where most prompts are (highly) unlikely to produce toxic output, but we still have an interest in understanding (and ultimately minimizing) the residual risk of toxicity in the model’s generated outputs.
@@
Figure 5: Frequency distribution of all words occurring in explanations of toxicity for LLama and OLMo.
@@
One potential application lies in red-teaming. Red-teaming is the practice by which organizations proactively (before deployment or on an ongoing basis) stress-test generative AI systems to uncover potential vulnerabilities, harmful behaviors, and/or failure modes (perez2022red).
@@ 5.3 Case Study 3: Sentiment classification in stories
For sentiment scoring, we apply the SentimentIntensityAnalyzer from NLTK (loper2002nltk), [^24] https://www.nltk.org/index.html which assigns a compound sentiment score to each generated output. The analysis focuses on the average sentiment score produced by the model across multiple (10) generations per prompt. As input the LLMs receive story prompts from a publicly available story generation dataset. [^25] https://huggingface.co/datasets/qwedsacf/story-generation For the case study, we consider prompts that produce negative sentiment on average. Thus the explanations reveal segments of the input without which the LLMs would not have produced negative sentiment.
@@
The complete results can be found in Table LABEL:tab:cf_results_sentiment_complete; we show an example in Table LABEL:tab:cf_results_sentiment_example. We find that in many cases the negative sentiment score of the output is maximal (1.00) or almost (e.g., 0.96). This can often be reduced to zero by removing only one or two key sentences. In a small number of cases, the algorithm did not produce an explanation within the allotted time limit. Since there are no overlapping sentences across the different stories (apart from the instruction prompt “Finish the following story.”), we do not present distribution plots.
@@
Once the “offending” sentences (or other elements) have been identified, they can be replaced, either manually or automatically, for example with paraphrased sentences. To illustrate, we compare two approaches to modify the prompt based on paraphrasing: (a) paraphrasing the sentences identified in the explanations, and (b) paraphrasing the same number of sentences selected at random. The original prompts yield an average negative sentiment score of 0.39. When sentences are paraphrased at random, the average sentiment decreases to 0.31, while paraphrasing based on the explanations further reduces it to 0.28.262626We use the following prompt to paraphrase the sentences with GPT-4o-mini: “You are a helpful assistant. Paraphrase the following sentence. Maintain the same meaning but use different wording:
@@ 6 Discussion and Conclusion
Another important but underexplored scenario is when LLMs are used directly as decision-makers. For example, in zero-shot or few-shot setups, LLMs may implicitly classify inputs—screening resumes, scoring essays, or triaging documents—without a separate classifier head. In these cases, providing counterfactual explanations for the final output may be essential for fairness auditing and regulatory compliance (wilson2024gender).
@@
The PCEs we present are correct—they indeed show elements of the input that cause the LLM to produce outputs with the focal characteristics. Whether they are useful in context depends on the reason(s) for producing them. As martens2025beware argue, explanation quality ultimately is context-dependent: what counts as a “good” explanation varies across stakeholders and across use cases. Explanations can be poor in a particular context because they are irrelevant, misleading or even harmful. It is important to separate out the production of explanations that are technically correct—the subject of this paper—with explanations that are good for a particular purpose and stakeholder. We think that establishing the former is necessary for the latter.
@@
The study of the utility of PCEs could be a significant next chapter in the decades-long tradition of research on explanations for AI systems.272727martens2014explaining provide a comprehensive review of the prior literature in IS and beyond. There seems to be an entire field of research waiting to be plowed, tilled, and planted. Once we can understand better the relationship between the input to generative AI systems and the output that they produce, how do we provide explanations that go further and are truly useful for different purposes? We suggested some different uses in the case studies, specifically prompt engineering and red teaming. Looking at prior work across XAI we see many more possibilities, especially once AI systems are not just producing output, but are taking actions that might affect customers and citizens. No single paper could possibly do justice to the topic; we need a concerted effort to spur research in the area generally and broadly. For example, the field would benefit from user studies focusing on whether generative AI explanations (such as PCEs) satisfy the desiderata of AI explanations laid out by theory (martens2014explaining).
@@
A well-known limitation is that counterfactual explanations are not well suited to understanding the default behavior of a system. This applies for generative AI as well as predictive AI. For example, in the context of political leaning, earlier research has shown that many LLMs inherently exhibit political leanings, producing responses that consistently skew left or right regardless of the input (bang2024measuring; buyl2024large). In this setting, PCEs will not necessarily be able to identify meaningful prompt-level explanations, as no change may have an effect on the focal characteristic. The PCE algorithm would simply fail to find any explanations. Despite this limitation, note that this outcome could be informative in itself: it suggests that the source of the political bias in the output lies within the underlying system, and not from the phrasing of the input.
@@
A different sort of limitation stems from the fact that generating PCEs can be very computationally expensive for large prompts or large sets of prompts, since it requires running the model $n$ times for each possible masking. Future work could address this inefficiency. One avenue for addressing this is to structure the search to be more efficient, for example by arranging the prompt elements hierarchically—for example, chapters, sections, sentences, words. Another promising direction is the use of diffusion-based language models (diffusion-LMs) (li2022diffusion), which allow the classifier to be consulted earlier in the generative process, thereby eliminating the need for a full generation phase and substantially reducing computational cost.
@@
Finally, our framework was developed and tested in the context of text-based language models. In principle, the approach could be extended to other generative modalities such as image or audio generation, opening up a broad range of future research directions.

### v2

@@ 4.1 Explanation algorithm
In the initial scoring phase, the PCE-1 algorithm evaluates the influence of candidate explanations (elements in the input prompt) on a specific output characteristic such as toxicity or sentiment. First, the original prompt $S$ is passed to a generative AI system $G$ to produce a number of outputs (with the number equal to num_samplesnum\_samples), presuming G is non-deterministic. These generated outputs are evaluated using downstream classifier $C_{m}$, which returns a score $p$ indicating how strongly the output exhibits the focal target property. As explained above, how this score is calculated depends on the downstream tasks. Next, if this score exceeds the predefined thresholdthreshold, the algorithm proceeds to an element-level sensitivity analysis for that prompt. For each element $e_{i}$ in the prompt (with in total $n$ elements), a new version of the prompt is created with that element replaced (e.g., masked) ($S\setminus e_{i}$). The modified prompt is fed again into the generative model, and the resulting outputs are scored by the classifier. These element-level scores are recorded in a dictionary, capturing how the removal of each element affects the output’s classification score.
@@
This explanation algorithm is intended as a first proof of concept for generating PCEs for a non-deterministic, generative AI system. Our default search strategy is an element-level textual masking approach, where we iteratively replace (“mask”) elements from the input with underscores ($_{}^{\prime}\_^{\prime}$) to test their causal effect on the classifier’s output. However, more sophisticated techniques could be explored, such as using system-specific masking tokens. [^15] As noted above, our experiments using system masking tokens yielded comparable results to using underscores for masking. Underscores are more broadly usable. As an alternative to masking, replacing input segments with alternative text may yield more natural and actionable counterfactuals (e.g, Kim et al. (2020)). In settings with very large input prompts, another alternative would be grouping phrases or semantically related elements and evaluating their collective influence as an initial filtering step could substantially improve the efficiency of the search process.
This approach could be applied to any large language model. For the experiments in this paper, we apply it to LLaMA 3.1–8B (Grattafiori et al., 2024) (hereafter LLaMA) and OLMo-2-0425-1B (Walsh et al., 2025) (hereafter OLMo). [^16] https://huggingface.co/meta-llama/Llama-3.1-8B, [^17] https://huggingface.co/allenai/OLMo-2-0425-1B We use the following parameter settings: num_samplesnum\_samples = 10 (except for toxicity, where we use num_samplesnum\_samples = 100) and num_explanationsnum\_explanations = 5, and time_limittime\_limit = 60 seconds. The returned numbers of explanations for each domain and model are listed in Table 1.
@@ 5 Case studies
We illustrate the production of PCEs for generative (LLM-based) systems with three case studies: [^18] Similarly to the prior work of Fernández-Loría et al. (2022). detection of political leaning, toxicity prediction, and sentiment classification. Each case study shows the explanations produced with case-specific settings for the algorithm parameters, and discusses case-specific implications. The case studies also demonstrate the versatility of counterfactual explanations across tasks with very different data distributions: in political leaning detection, the distribution is bimodal; in toxicity detection, the signal is concentrated in the relatively few toxic instances; and in sentiment analysis, the focus is on the long tail.
@@ 5.1 Case Study 1: Detection of political leaning
LLM-based AI systems have been criticized for producing biased outputs (Gallegos et al., 2024; Ho et al., 2025; Goethals et al., 2026). One type of bias that is often examined is when the generated outputs systematically lean toward one side of the political spectrum (Motoki et al., 2024; Rettenberger et al., 2025; Anthropic, 2025). Stakeholders investing in, building, and integrating AI systems are interested in (a) whether their AI systems’ generations do in fact lean one direction or the other politically, (b) if so, are there specific things in the prompts that cause the systems to generate such biased outputs? And (c) what they might do about it. Prior work has created classifiers to address (a), which we discuss below. This paper presents a solution to (b). A full answer to part (c) is beyond the scope of this paper (and is very complex); we will provide some ideas for (c) toward the end of this section. [^19] OpenAI recently reported on their desire that their LLM-driven AI systems not exhibit political bias, as well as what they are doing about it, stating clearly that “ChatGPT shouldn’t have political bias in any direction.” They estimate that in live traffic “less than 0.01% of all ChatGPT responses show any signs of political bias.” We note that this still would result in a non-negligible number of biased responses, given the tremendous volume of queries to ChatGPT. https://openai.com/index/defining-and-evaluating-political-bias-in-llms/. Similarly, Anthropic announced in a blog post that they are reducing and actively monitoring political bias (Anthropic, 2025).
@@
Generating news articles from headlines is a well-established method for evaluating political bias in language models (Bang et al., 2024). To assess the political leaning of the generated outputs, we rely on PoliticalBiasBERT, a classifier designed to predict the dominant political ideology of news content that is widely used (Bucket Research, 2023; Baly et al., 2020). This model produces as output one of three categorical labels—left, center, or right. As noted above, we apply PoliticalBiasBERT not to the headlines, but to the generated output—in line with the architecture in Figure 1. Thus the headline is the prompt and the news article is the generated output, which aligns with many uses of LLM-based systems, taking a relatively short prompt and producing a significantly longer output.
@@
As discussed in the literature on counterfactual explanation (Martens and Provost, 2014; Wachter et al., 2017; Verma et al., 2024), there are many goals for producing such explanations. One common goal is to help managers and developers understand the reasons for undesirable system behavior. These results illustrate how PCEs can inform these stakeholders about the elements of the input prompts that led to the political leaning.
@@
We see from the figure that there are indeed words that consistently produce right-leaning output. Such an analysis could be helpful to data scientists, developers, and domain experts working to improve the AI model; a common argument for including counterfactual explanations in the model-development toolkit (Martens and Provost, 2014; Abid et al., 2022; Gan et al., 2021; Yousefzadeh and O’Leary, 2019). Specifically, the development process can focus on understanding why the presence of these particular words in prompts leads to the biased output, and gives a focused target for reducing the bias.
@@
Deciding what to do based on what counterfactual explanations reveal depends not only on the problem setting but also on the stakeholder in question (Langer et al., 2021). A top-notch AI engineer would have more options at her disposal than a business developer working on an LLM-driven solution. One potential broad application of PCEs for generative outputs is prompt engineering. Knowing specifically what caused a desired or undesired output characteristic could help to engineer future prompts that produce or avoid desired or undesired generations. A comprehensive investigation of prompt engineering driven by PCEs is well beyond the scope of this paper (and would be a significant contribution as future work). As a follow-up for this case study, we demonstrate the potential as follows.
The generalizability of the content of the explanations suggests that the PCEs could inform prompt engineering strategies for avoiding the focal characteristic (here, right-leaning) in the LLM generations. To illustrate, we conducted the following comparison. We (a) again selected the 100 headlines that contained the highest frequency of the top 20 politically salient words, and (b) created a corresponding set of headlines in which we replaced these top 20 words with synonyms (e.g., replacing “lawsuit” with “legal case” and “Netanyahu” with “Israeli prime minister”). [^21] We asked ChatGPT to provide synonyms for each of the words The results are shown in Table 4.
@@ 5.2 Case Study 2: Toxicity prediction
LLMs carry a well-documented risk of producing toxic or otherwise objectionable content. This issue to a large extent arises from the nature of their training data—large-scale internet corpora that inevitably include hate speech, stereotypes, and offensive language (Abid et al., 2021; Bender et al., 2021; Gehman et al., 2020). As a result, even when prompted with seemingly neutral inputs, LLMs can generate responses that violate corporate policies and social or content-safety norms. While these risks remain a concern, they are substantially mitigated through extensive post-training interventions implemented by the providers. Nonetheless, although rare, the systems still sometimes generate toxic outputs. The consequences of such outputs range from diminished trust in AI systems to reputational and legal risks for the organizations deploying them (Weidinger et al., 2022). Thus both LLM providers and organizations integrating them have an interest in understanding and avoiding toxic generations.
To study the relationship between prompts and toxic generated outputs, we use the RealToxicityPrompts dataset (Gehman et al., 2020), which contains a mix of prompts that either are neutral or were known to induce toxic outputs under some conditions. [^22] https://huggingface.co/datasets/allenai/real-toxicity-prompts Although the vast majority of prompts are unlikely to yield toxic responses from current LLMs, we observe that toxicity indeed still emerges. To determine whether a given output is toxic, we use the RoBERTa-based toxicity classifier developed by Logacheva et al. (2022), which outputs a binary label for toxicity. [^23] https://huggingface.co/s-nlp/roberta_toxicity_classifier This case study examines the setting where most prompts are (highly) unlikely to produce toxic output, but we still have an interest in understanding (and ultimately minimizing) the residual risk of toxicity in the model’s generated outputs.
@@
Figure 5: Frequency distribution of all words occurring in explanations of toxicity for LLama and OLMo. Most words appear in only one explanation, but in each case we see a dozen or so words that occur across multiple explanations.
@@
One potential application lies in red-teaming. Red-teaming is the practice by which organizations proactively (before deployment or on an ongoing basis) stress-test generative AI systems to uncover potential vulnerabilities, harmful behaviors, and/or failure modes (Perez et al., 2022).
@@ 5.3 Case Study 3: Sentiment classification in stories
For sentiment scoring, we apply the SentimentIntensityAnalyzer from NLTK (Loper and Bird, 2002), [^26] https://www.nltk.org/index.html which assigns a compound sentiment score to each generated output. The analysis focuses on the average sentiment score produced by the model across multiple (10) generations per prompt. As input the LLMs receive story prompts from a publicly available story generation dataset. [^27] https://huggingface.co/datasets/qwedsacf/story-generation For the case study, we consider prompts that produce negative sentiment on average. Thus the explanations reveal segments of the input without which the LLMs would not have produced negative sentiment.
@@
The complete results can be found in Table LABEL:tab:cf_results_sentiment_complete; we show an example in Table LABEL:tab:cf_results_sentiment_example. We find that in many cases the negative sentiment score of the output initially is maximal (1.00) or almost (e.g., 0.96). This can often be reduced to zero by masking only one or two key sentences. In a small number of cases, the algorithm did not produce an explanation within the allotted time limit. Since there are no overlapping sentences across the different stories (apart from the instruction prompt “Finish the following story.”), we do not present distribution plots.
@@
Once the “offending” sentences (or other elements) have been identified, they can be replaced, either manually or automatically, for example with paraphrased sentences. To illustrate, we compare two approaches to modify the prompt based on paraphrasing: (a) paraphrasing the sentences identified in the explanations, and (b) paraphrasing the same number of sentences selected at random. The original prompts yield an average negative sentiment score of 0.39. When sentences are paraphrased at random, the average sentiment decreases to 0.31, while paraphrasing based on the explanations further reduces it to 0.28.282828We use the following prompt to paraphrase the sentences with GPT-4o-mini: “You are a helpful assistant. Paraphrase the following sentence. Maintain the same meaning but use different wording: An example of a use case can be found in Meske and Bunde (2023). They demonstrate how using a user interface with local explanations in hate speech detection leads to users experiencing an increased understanding of the work processes involved in the artifact and increased trustworthiness compared to an interface without local explanations.
@@ 6 Discussion and Conclusion
Another important but underexplored scenario is when LLMs are used directly as decision-makers. For example, in zero-shot or few-shot setups, LLMs may implicitly classify inputs—screening resumes, scoring essays, or triaging documents—without a separate classifier head. In these cases, providing counterfactual explanations for the final output may be essential for fairness auditing and regulatory compliance (Wilson and Caliskan, 2024).
@@
The PCEs we present are correct—they indeed show elements of the input that cause the LLM to produce outputs with the focal characteristics. Whether they are useful in context depends on the reason(s) for producing them. As Martens et al. (2025) argue, explanation quality ultimately is context-dependent: what counts as a “good” explanation varies across stakeholders and across use cases. Explanations can be poor in a particular context because they are irrelevant, misleading or even harmful. Bauer et al. (2023) for example highlight how AI explanations lead to mental model adjustments which could manipulate user behavior. It is important to separate out the production of explanations that are technically correct—the subject of this paper—with explanations that are good for a particular purpose and stakeholder. We think that establishing the former is necessary for the latter.
@@
The study of the utility of PCEs could be a significant next chapter in the decades-long tradition of research on explanations for AI systems. [^29] Martens and Provost (2014) provide a comprehensive review of the prior literature in IS and beyond. There seems to be an entire field of research waiting to be plowed, tilled, and planted. Once we can understand better the relationship between the input to generative AI systems and the output that they produce, how do we provide explanations that go further and are truly useful for different purposes? We suggested some different uses in the case studies, specifically prompt engineering and red teaming. Looking at prior work across XAI we see many more possibilities, especially once AI systems are not just producing output, but are taking actions that might affect customers and citizens. No single paper could possibly do justice to the topic; we need a concerted effort to spur research in the area generally and broadly. For example, the field would benefit from user studies focusing on whether generative AI explanations (such as PCEs) satisfy the desiderata of AI explanations laid out by theory (Martens and Provost, 2014).
@@
A well-known limitation is that counterfactual explanations are not well suited to understanding the default behavior of a system. This applies for generative AI as well as predictive AI. For example, in the context of political leaning, earlier research has shown that many LLMs inherently exhibit political leanings, producing responses that consistently skew left or right regardless of the input (Bang et al., 2024; Buyl et al., 2024). In this setting, PCEs will not necessarily be able to identify meaningful prompt-level explanations, as no change may have an effect on the focal characteristic. The PCE algorithm would simply fail to find any explanations. Despite this limitation, note that this outcome could be informative in itself: it suggests that the source of the political bias in the output lies within the underlying system, and not from the phrasing of the input.
@@
A different sort of limitation stems from the fact that generating PCEs can be very computationally expensive for large prompts or large sets of prompts, since it requires running the model $n$ times for each possible masking. Future work could address this inefficiency. One avenue for addressing this is to structure the search to be more efficient, for example by arranging the prompt elements hierarchically—for example, chapters, sections, sentences, words. Another promising direction is the use of diffusion-based language models (diffusion-LMs) (Li et al., 2022), which allow the classifier to be consulted earlier in the generative process, thereby eliminating the need for a full generation phase and substantially reducing computational cost.
@@
Finally, our framework was developed and tested in the context of text-based language models. In principle, the approach could be extended to other generative modalities such as image (similar to Meske and Bunde (2020) and to Vermeire et al. (2022)) or audio generation, opening up a broad range of future research directions.